OpenStudy (anonymous):

x^3-2x^2-x+2=0 how to find x?

OpenStudy (math&ing001):

You could factor it this way: x^2(x-2) - (x-2) = (x^2-1)(x-2) = 0 Should be easier to solve now.

OpenStudy (anonymous):

or use the rational root theorem